# SHIELD: Suppressing Hallucinations In LVLM Encoders via Bias and VuLnerability Defense

### v1

@@ A.1 GPT4o Assisted Evaluation
Appendix B More Results
B.1 POPE Evaluation on AOKVQA & GQA
@@ B.2 Detailed Results on MME
Table 9 presents the results on perception tasks of the MME benchmark using the LLaVA-1.5 7B model. Compared to the vanilla LVLM, SHIELD achieves overall improvements, highlighting its ability to reduce hallucinations and enhance perception capabilities. This improvement likely stems from SHIELD’s effectiveness in mitigating biases and vulnerabilities, thereby recalibrating the LVLM’s visual feature extraction. When compared to previous methods, SHIELD achieves a higher total perception score (Table 6) but exhibits limited improvements on tasks requiring external knowledge beyond the input image (e.g., Celebrity, Scene, Landmark, Artwork). This limitation may result from contrastive decoding in SHIELD, which directs the LVLM to prioritize visual inputs over leveraging prior knowledge embedded in the LLM.
Furthermore, Table 10 presents the results on cognition tasks of the MME benchmark using the LLaVA-1.5 7B model. The findings indicate that applying SHIELD improves the LVLM’s recognition on complex visual scenes, such as common sense reasoning, but performs poorly on simpler visual scenes, including numerical calculation, text translation, and code reasoning. This may be because simpler visual inputs are less likely to induce hallucinations, while contrastive decoding in SHIELD may limit the utilization of prior knowledge embedded in the LLM.
Table 11: Inference Overhead on POPE COCO using LLaVA-1.5
Method Accuracy $\uparrow$ F1 $\uparrow$ Time/Sample (ms) $\downarrow$ Relative $\downarrow$
Vanilla 81.3 79.6 128 1.0×
VCD 84.6 84.4 190 1.5×
OPERA 84.7 85.4 1586 12.4×
@@
B.3 Inference Overhead
Table 11 reports the per-sample inference time on the POPE COCO subset with the LLaVA-1.5 7B model. SHIELD introduces more overhead than lightweight methods such as VCD but remains faster than OPERA. Considering its strong reduction of hallucinations, this trade-off is reasonable. The overhead can be further alleviated by shortening naive captions or reducing adversarial steps.
@@ Appendix C Additional Hyper-Parameters Ablation
C.1 $\alpha$ Ablation
Table 15 summarizes the ablation results for the parameter $\alpha$, which regulates the influence of contrastive decoding when integrated with adversarial attack strategies. The study shows a significant reduction in hallucinations on the CHAIR benchmark as $\alpha$ increases from 1.0 to 2.0, demonstrating the effectiveness of addressing vulnerabilities.
C.2 $\beta$ Ablation
@@
Table 15 reports the ablation results for the parameter $\beta$, which governs the adaptive plausibility constraint. A higher value of $\beta$ enforces stronger truncation, preserving only tokens with high generation probabilities. As $\beta$ increases, changes in hallucination reduction exhibit minor fluctuations, suggesting that while the adaptive plausibility constraint, as an integral part of contrastive decoding, prevents the generation of implausible content, it plays a limited role in alleviating object hallucinations.
C.3 $K$ Ablation
@@
Table 15 reports the ablation results for the parameter $K$, which determines the number of noise samples employed to estimate and subsequently eliminate inherent bias from the visual tokens. Increasing $K$ improves the accuracy of inherent bias estimation, resulting in more effective hallucination mitigation. However, when $K$ becomes excessively large, the estimated bias converges toward zero, limiting its impact on mitigating hallucinations.
C.4 $l$ Ablation
@@
Table 15 reports the ablation results for the learning rate $l$ used in the vulnerability addressing, which regulates the granularity of adversarial attack tensor calculations. When the learning rate is too large, the generated attack tensor fails to adapt effectively to image details, reducing its effectiveness. Consequently, the subsequent contrastive decoding process cannot adequately minimize hallucinations caused by vulnerabilities in the visual encoder.
@@
Appendix D Case Study

### v2

@@ A.1 GPT-4o Assisted Evaluation
Appendix B More Results
B.1 POPE Evaluation on AOKVQA & GQA
@@ B.2 Detailed Results on MME
Table 12 presents the results on perception-related tasks of the MME benchmark using the LLaVA-1.5 7B model. Compared to the vanilla LVLM, SHIELD achieves overall improvements, highlighting its ability to reduce hallucinations and enhance perception capabilities. This improvement likely stems from SHIELD’s effectiveness in mitigating biases and vulnerabilities, thereby recalibrating the LVLM’s visual feature extraction. When compared to previous methods, SHIELD achieves a higher total perception score (Table 7) but exhibits limited improvements on tasks requiring external knowledge beyond the input image (e.g., Celebrity, Scene, Landmark, Artwork). This limitation may result from contrastive decoding in SHIELD, which directs the LVLM to prioritize visual inputs over leveraging prior knowledge embedded in the LLM.
Furthermore, Table 13 showcases the results on cognition-related tasks within the MME benchmark using the LLaVA-1.5 7B model. The results show that applying SHIELD improves the LVLM’s recognition on complex visual scenes (e.g., Common Sense Reasoning) but performs poorly on simpler visual scenes (e.g., Numerical Calculation, Text Translation, and Code Reasoning). This may be because simpler visual inputs are less likely to induce hallucinations, while contrastive decoding in SHIELD may limit the utilization of prior knowledge embedded in the LLM.
Table 14: GPT-4 Assisted Evaluation on MMHal-Bench
Method Overall $\uparrow$ Hallucination $\downarrow$ Attr. $\uparrow$ Adv. $\uparrow$ Comp. $\uparrow$ Cnt. $\uparrow$ Rel. $\uparrow$ Env. $\uparrow$ Hol. $\uparrow$ Other $\uparrow$
Vanilla 2.21 0.70 2.75 2.00 2.50 2.50 1.67 2.17 1.83 2.25
Ours 2.65 0.61 2.42 1.83 2.58 2.67 2.83 3.92 2.08 2.83
B.3 GPT-4 Assisted Evaluation on MMHal-Bench
Table 14 presents results on MMHal-Bench (Sun et al., 2024b), a benchmark designed to evaluate hallucinations in open-ended multimodal tasks. It spans eight hallucination-related categories, including attribute, adversarial object, comparison, counting, relation, environment, holistic description, and other visually grounded errors. To ensure up-to-date evaluation, we re-assess both the vanilla LLaVA-1.5 7B and our method using the latest version of GPT-4, as the original benchmark relied on the deprecated gpt-4-0314. Our method significantly reduces hallucination and improves performance across most categories, with the most notable gains observed in relation and environment understanding.
Table 15: Extended Evaluation on Recent VLMs Variants (POPE COCO Adversarial Split)
Method Qwen2-VL (7B) Qwen2.5-VL (7B) Qwen3-VL (8B)
Accuracy ($\uparrow$) F1 Score ($\uparrow$) Accuracy ($\uparrow$) F1 Score ($\uparrow$) Accuracy ($\uparrow$) F1 Score ($\uparrow$)
Vanilla 85.3 84.5 85.6 85.1 86.2 85.8
Ours 86.1 85.4 86.8 86.0 87.4 86.6
B.4 Extended Evaluation on Recent VLMs
The VLMs used in our main experiments (LLaVA-1.5, InstructBLIP, and Qwen-VL) follow standard practice in recent multimodal hallucination studies, ensuring fair and reproducible comparison. To further evaluate the generalization of SHIELD, we extend our experiments to more recent Qwen-VL variants: Qwen2-VL (7B), Qwen2.5-VL (7B), and the newly released Qwen3-VL (8B). Notably, Qwen3-VL 8B was released in November 2025 and represents the latest generation of the Qwen-VL family. We report results on the POPE COCO adversarial split.
As shown in Table 15, SHIELD consistently improves both accuracy and F1 score across all models. These results demonstrate the effectiveness of our approach on more capable and recently released VLMs.
Table 16: Comparison with Training-Free and Training-Based Methods on CHAIR (512 tokens)
Method Training $C_{S}$($\downarrow$) $C_{I}$($\downarrow$)
Vanilla 48.8 14.2
DoLa (Chuang et al., 2024) 47.7 13.8
ICD (Wang et al., 2024) 47.4 13.9
VCD (Leng et al., 2024) 46.8 13.2
TAME (Tang et al., 2025) 45.2 14.0
OPERA (Huang et al., 2024) 44.8 12.8
SID (Huo et al., 2025) 44.2 12.2
LLaVA-RLHF (Sun et al., 2024b) 43.6 10.5
CCA-LLaVA (Xing et al., 2024) 43.0 11.5
Less is More (Yue et al., 2024) 40.2 12.3
MCA-LLaVA (Zhao et al., 2025b) 38.0 10.9
@@ B.4 Extended Evaluation on Recent VLMs
B.5 Extended Comparison with Existing Hallucination Mitigation Methods
To provide a more comprehensive evaluation, we expand our comparison to include a wider set of hallucination mitigation methods on the CHAIR benchmark. These include both training-free methods (DoLa, ICD, VCD, TAME, OPERA, SID, and Ours) and training-required approaches (LLaVA-RLHF, CCA-LLaVA, Less is More, and MCA-LLaVA). All methods are implemented with LLaVA-1.5 7B to ensure consistency in backbone and evaluation setup.
As shown in Table 16, our method outperforms all training-free baselines and remains competitive with training-based methods. Notably, SHIELD achieves the lowest hallucination rates on both sentence and instance levels while requiring no additional VLM fine-tuning.
@@ Appendix C Additional Hyper-Parameters Ablation
C.1 Effect of $\alpha$ in Contrastive Decoding
Table 20 presents the results of an ablation study on $\alpha$, which controls the impact of contrastive decoding combined with adversarial attacks. The study shows a significant reduction in hallucinations on the CHAIR benchmark as $\alpha$ increases from 1.0 to 2.0, demonstrating the effectiveness of addressing vulnerabilities.
@@ C.1 Effect of $\alpha$ in Contrastive Decoding
C.2 Effect of $\beta$ in Adaptive Plausible Constraint
@@
Table 20 presents the results of an ablation study on $\beta$, which controls the adaptive plausibility constraint. A larger $\beta$ indicates more aggressive truncation, retaining only high-probability tokens. As $\beta$ increases, changes in hallucination reduction exhibit minor fluctuations, suggesting that while the adaptive plausibility constraint, as an integral part of contrastive decoding, prevents the generation of implausible content, it plays a limited role in alleviating object hallucinations.
C.3 Effect of $K$ in Reducing Inherent Bias
@@
Table 20 presents the results of an ablation study on $K$, which specifies the number of noise inputs used to estimate inherent bias for subsequent removal from visual tokens. Increasing $K$ improves the accuracy of inherent bias estimation, resulting in more effective hallucination mitigation. However, when $K$ becomes excessively large, the estimated bias converges toward zero, limiting its impact on mitigating hallucinations.
C.4 Effect of $l$ in Addressing Vulnerability
@@
Table 20 presents the results of an ablation study on the learning rate within Vulnerability Addressing, which controls the granularity of adversarial attack tensor computation. When the learning rate is too large, the generated attack tensor fails to adapt effectively to image details, reducing its effectiveness. Consequently, the subsequent contrastive decoding process cannot adequately minimize hallucinations caused by vulnerabilities in the visual encoder.
Table 21: Attack Steps Ablation
Steps $C_{S}\downarrow$ $C_{I}\downarrow$
15 40.8 11.6
30 36.6 10.3
45 36.6 10.4
C.5 Effect of Attack Step in Addressing Vulnerability
Table 21 presents an ablation study on the number of adversarial optimization steps in addressing vulnerability module. This parameter controls how thoroughly the attack tensor explores encoder weaknesses. The results show that performance saturates at 30 steps, with no further gains at 45 steps.
Appendix D Additional Analysis
Table 22: Class-wise hallucination rate (%) under different noise settings on POPE COCO
Input person skis snowboard bird backpack skateboard bowl scissors knife keyboard carrot refrigerator remote sandwich sink
Noise Img 100 80 30 30 20 20 20 10 10 10 10 10 10 10 0
Noise Tok 0 10 10 0 10 0 0 10 0 10 0 0 10 0 20
D.1 Additional Inherent Bias Analysis
To further examine inherent bias in the visual encoder, we conduct a controlled experiment using the POPE COCO subset. For each object category, we design 10 identical POPE-style questions (e.g., “Is there a {object} in the image?”) and evaluate hallucination rates under two distinct noise settings. In the first setting, the visual encoder receives noise as image input (Noise Image), while in the second, the original image is retained but its visual tokens are replaced with noise embeddings (Noise Visual Tokens).
We measure the proportion of “Yes” responses, which indicate hallucination, for each object. As shown in Table 22, hallucinations under the Noise Image setting are heavily skewed, with certain categories (e.g., person) being disproportionately predicted. In contrast, hallucination rates under Noise Visual Tokens remain uniformly across categories. These findings confirm that the visual encoder exhibits class-specific inherent biases, which can be effectively revealed through noise-based evaluation.
To mitigate the impact of randomness from individual noise inputs, SHIELD averages responses over $K$ noise inputs. This aggregation helps reduce variance and provides a more stable estimate of the encoder’s bias tendencies.
Figure 9: Example naive captions with and without plausible hallucination. The hallucinated version includes a refrigerator not present in the image.
Table 23: Impact of Plausible Hallucinations in Naive Caption
Model Accuracy ($\uparrow$) F1-Score ($\uparrow$)
Vanilla 78.90 77.50
Ours w/ Raw Caption 80.13 79.14
Ours w/ Hallucinated Caption 79.83 78.81
D.2 Plausible Hallucination in Naive Caption Analysis
To assess the effect of plausible hallucinations in naive captions, we conduct a controlled experiment on the POPE COCO adversarial subset. We simulate plausible hallucinations by introducing mild noise into the visual input during naive caption generation. This can produce hallucinated objects that are contextually reasonable but not actually present in the image.
Figure 9 shows one representative example where the hallucinated caption mentions a “refrigerator”, which is plausible in the context but does not appear in the image. Despite this, grounded elements such as “microwave” and “sign” remain in the caption, which continue to support effective token re-weighting. The impact of such plausible hallucinations is limited because they typically have lower visual grounding and do not dominate the token distribution.
To quantify this, we compare SHIELD’s performance when using raw versus hallucinated captions. As shown in Table 23, the drop in performance is minor, indicating robustness of SHIELD to plausible caption noise.
Table 24: Effect of Attack Strategy on CHAIR
Configuration $C_{S}$ ($\downarrow$) $C_{I}$ ($\downarrow$)
Vanilla 48.8 14.2
Ours w/ FGSM-based Vulnerability Module 39.2 11.5
Ours w/ PGD-based Vulnerability Module 37.2 11.3
Ours (Learnable Attack) 36.6 10.3
D.3 Attack Strategy Analysis
To evaluate the design of our vulnerability addressing module, we compare our learnable perturbation with standard adversarial attacks that use fixed step sizes, such as FGSM (Goodfellow et al., 2015) and PGD (Madry et al., 2018). Instead of relying on fixed-magnitude updates, our method learns a fine-grained perturbation through adversarial optimization between the visual encoder and its paired CLIP text encoder (see Section 3.2.4). Conditioned on both the image and naive caption, the perturbation is specifically optimized to expose vulnerabilities in the encoder, offering more flexible and precise control than fixed-step-size attacks.
To validate its effectiveness, we evaluate multiple configurations on the CHAIR benchmark, including variants where the vulnerability addressing module is implemented with FGSM or PGD instead. As shown in Table 24, while all configurations incorporating the vulnerability module help reduce hallucinations, our learnable attack achieves the lowest scores on both $C_{S}$ and $C_{I}$.
@@ D.3 Attack Strategy Analysis
Appendix E Case Study
@@
Appendix F The Use of Large Language Models (LLMs)
We used GPT for two purposes: (i) polishing grammar and improving readability, and (ii) assisting in the evaluation of LVLM outputs. All research ideas and analyses were conducted by the authors, who take full responsibility for the content.